Circular No. 7888

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002cq IN ANONYMOUS GALAXY W. M. Wood-Vasey, G. Aldering, P. Nugent, and R. Rudas, Lawrence

Berkeley National Laboratory, report the discovery of an apparent supernova (mag 19.2) on an unfiltered image taken on Apr. 20.3 UT with the 1.2-m Schmidt telescope at Palomar in a joint collaboration with the NEAT program (E. F. Helin, S. Pravdo, M. Hicks, and K. Lawrence, Jet Propulsion Laboratory). The object retained the same brightness on additional images taken on Apr. 29.2. SN 2002cq is located at α = 12^{h}38^{m}40^{s}.43, δ = −2^{o}58′13′′.4 (equinox J2000.0), which is 15′′ south and 10′′ west of the apparent host galaxy. Nothing is visible at this location on images from Mar. 21 and Apr. 3 (limiting mag 20).

XTE J0929–314 R. A. Remillard, Massachusetts Institute of Technology (MIT), for

the RXTE ASM Team at MIT and Goddard Space Flight Center, reports the discovery of a faint x-ray transient at α = 9^{h}29^{m}22^{s}, δ = −31^{o}22′.8 (equinox J2000.0; estimated uncertainty 3′, 90-percent confidence). The source is visible in sky maps computed from 6-day intervals of ASM data. Average fluxes (2–12 keV): Apr. 13–18, 15 ± 2 mCrab; Apr. 19–24, 20 ± 2; Apr. 25–30, 26 ± 1. The error circle contains no noteworthy sources in the Simbad catalogues. The ASM hardness ratios suggest an x-ray spectrum similar to transients associated with weakly magnetic neutron stars or some blackhole systems. Optical and radio observations are encouraged.

COMET C/2002 F1 (UTSUNOMIYA)

Ephemeris extension to IAUC 7869 (elements MPC 45335):

2002TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Apr. 24 2^{h}36.^{m}94 +31^{o}53.^{′}0 1.205 0.439 20.^{o}5 53.^{o}2 5.3
28 3 18.08 +29 12.7 1.243 0.458 20.2 49.4 5.6
May 2 3 52.09 +25 35.3 1.291 0.497 20.6 45.5 6.0
6 4 19.59 +21 37.5 1.346 0.551 21.5 42.2 6.6
10 4 42.02 +17 41.4 1.406 0.614 22.7 39.4 7.1
14 5 00.73 +13 57.8 1.468 0.682 24.0 37.1 7.7
18 5 16.74 +10 30.2 1.531 0.754 25.4 35.1 8.2
22 5 30.79 + 7 18.9 1.594 0.827 26.7 33.4 8.7
26 5 43.37 + 4 22.7 1.658 0.900 28.1 32.0 9.1
2002 April 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green